(a) Discuss the various price indices used to measure inflation in Pakistan.

(b) Choose an article on inflation from any Economics/ Business newspaper or magazine and discuss the current situation of inflation in Pakistan. (Don’t forget to mention the Publication and date)


1
Expert's answer
2021-04-15T20:42:41-0400
"solution"

A price index (PI) is a measure of how prices change over a period of time, or in other words, it is a way to measure inflation. There are multiple methods on how to calculate inflation.

[1]THE LASPEYRES FORMULA

The Laspeyres formula is generally used. for example:

Pakistan inflation rate for 2019 was 10.58%, a 5.5% increase from 2018.

The formula for Laspeyres Price Index is as follows:


 "L(P)=\frac{\Sigma_{i=1}^{i=N}P_i^FQ_i^B}{\Sigma_{i=1}^{i=N}P_i^BQ_i^B}"


 Where:

  • "P_i^\ B" : The price of good"\ i" in the Base period
  • "P_i^\ F" : The price of good"\ i" in the Final period
  • "Q_i^\ B" : The quantity consumed of good"\ i" in the Base period
  • "Q_i ^\ F" : The quantity consumed of good"\ i" in the Final period

 

The Laspeyres PI weighs prices (both Base period prices and Final period prices) with base period quantities. Consider an economy with N goods and services. The numerator in the Laspeyres price index calculates nominal expenditure required to consume base period quantity at final period prices. The denominator calculates nominal GDP in the base period.

 

[2]FISHERS PRICE INDEX.

Irving Fisher (1867-1947) was an American economist and statistician. Fisher was one of the earliest neoclassical economists in the US and is known as the first econometrician (application of linear regression to economic theory). The American worked on many areas of economics, including trade, monetary theory, and inflation measurement.

The formula for the Fisher Price Index is as follows:


 "F(P)=\sqrt{L(P)P(P)}"


The Fisher Price Index is the geometric mean of the Laspeyres Price Index and the Paasche Price Index.

[3]PAASCHE PRICE INDEX

Hermann Paasche (1851-1925) was a German economist and statistician. Paasche’s main contribution to economics and statistics was his work on wage inflation.

The formula for the Paasche Price Index is as follows:


 "P(P)=\frac{\Sigma_{i=1}^{i=N}P_i^BQ_i^F}{\Sigma_{i=1}^{i=N}P_i^BQ_i^F}"

 Where:

  • "P_i^B" : The price of good"\ i" in the Base period
  • "P_i^F" : The price of good"\ i" in the Final period
  • "Q_i^B" : The quantity consumed of good"\ i" in the Base period
  • "Q_i^F" : The quantity consumed of good"\ i" in the Final period

 

Paasche PI weighs prices (both Base period prices and Final period prices) with Final period quantities. Consider an economy with N goods and services. The numerator calculates nominal GDP in the Final period. The denominator in the price index calculates nominal expenditure required to consume Final period quantities at Base period prices.

 

[4]MARSHALL-EDGEWORTH INDEX

Alfred Marshall (1842-1924) was an English economist who is widely considered to be the father of modern neoclassical economics. Marshall’s book, Principles of Economics (1890), is one of the most influential textbooks in the history of economic thought. Francis Ysidro Edgeworth (1845-1926) was an Anglo-Irish economist and philosopher. Edgeworth was one of the earliest proponents of using statistics to analyze economic questions.

The formula for the Marshall-Edgeworth Price Index is as follows:

 

"ME(P)=\frac{L(P)+P(P)}{2}"


Where:

  • "L(P)" : The Laspeyres Price Index
  • "P(P)" : The Paasche Price Index

 

The Marshall-Edgeworth Price Index is the arithmetic mean (simple average) of the Laspeyres Price Index and the Paasche Price Index.


[B]

Inflation is once again the news, having risen further to 8.2 per cent in February. That inflation comes with costs is not news. For the poor, a rise in the prices of essential items (if it exceeds income growth) can be a death knell, both literally (for subsistence households), and indirectly, due to the inability to afford needed medical and health spending. It can also force parents to choose between whether their child goes to school or works.

Insofar as inflation erodes trust in the national currency as a store of value, it also erodes the associated national pride, and this is felt by all citizens. In Pakistan, this erosion has been significant: by the mid-1970s, the Pakistani rupee had lost half of the purchasing power it had in 1956; and by the early 1990s, it had lost 90pc

Moderate inflation, in the 3pc to 6pc range is generally considered desirable, and inflation below 3pc can actually be risky. Why? Moderate inflation can serve as a useful signal of demand pressures in normal times, and also lends flexibility to an economy adjusting to adverse shocks: if inflation is near zero, disinflation must involve nominal wage cuts, which are politically difficult.

For much of Pakistan’s history, inflation has been moderate, with two noticeable exceptions: 1972-76 and 2008-14, both of which coincided with record-high international oil prices; and followed/ accompanied public or private spending booms. Although inflation has picked up in the past few months and is now in the upper single digits, its level is still low by recent historical standards.

In sum, inflation is a multi-source problem. It has been high, but manageable, in Pakistan. But because it affects the poor disproportionately, the government must continue to take structural measures to keep it low, and to compensate the poor via lifeline tariffs and cash transfers for any temporary surges.


Published in Dawn, March 11th, 2020